A.2 Price Restrictions Question 4: Suppose demand and supply for a commodity are Q2P 11 and Q 2P 8 a. For that commodity, compute equilibrium price, b. Suppose that commodity with Q211 and Q 2P 8 were c. Now instead of labor, suppose that commodity with Qd-2P11 quantity, consumer surplus, and producer surplus. unskilled labor and a minimum wage of $5 is imposed. Compute the effect of that minimum wage on the unskilled labor. and Q 2P 8 were medical care and a maximum price of $4 is imposed to make medical care more affordable. Compute the effect of that maximum price on those seeking medical care. Explain your answers. Answer to Question:
Answer #1

a. Equilibrium price
Qd=Qs
-2P+11=2P-8
19=4P
P = 4.75
Qd= -2P+11=-2*4.75+11 = 1.5
CS= 0.5*1.5*(5.5-4.75)= 0.5625
PS=0.5*1.5*(4.75-4)= 0.5625

P

Qd

Qs

0.00

11.00

-8.00

0.25

10.50

-7.50

0.50

10.00

-7.00

0.75

9.50

-6.50

1.00

9.00

-6.00

1.25

8.50

-5.50

1.50

8.00

-5.00

1.75

7.50

-4.50

2.00

7.00

-4.00

2.25

6.50

-3.50

2.50

6.00

-3.00

2.75

5.50

-2.50

3.00

5.00

-2.00

3.25

4.50

-1.50

3.50

4.00

-1.00

3.75

3.50

-0.50

4.00

3.00

0.00

4.25

2.50

0.50

4.50

2.00

1.00

4.75

1.50

1.50

5.00

1.00

2.00

5.25

0.50

2.50

5.50

0.00

3.00

6.00 5.00 4.50 4.00 Price 3.00 2.50 2.00 1.50 1.00 0.50 0100 9 87 -6 -5 -4 -3 2 -1 0 1 23 4 5 6 7 8 9 10 11 12 Quantity

b. With a minimum wage of $5, there would be excess supply than demand leading to surplus labor and causing unemployment.

6.06Unemployment 5.00 4.50 4.00 3.00 2.00 1.50 0.50 0.00ーーー ㄧ ㄧㄒㄧ- ㄒ ㄧㄧ ㄧㄒ 9 8 7 -6 -5 -4 -3 2 -1 0 1 23 4 5 6 7 8 9 10 11 12

c. The effect of maximum price of $4 is zero supply, short supply but the excess demand of 3 quantity, consumer surplus is 0.5*1.5*3= 2.25, producer surplus = 0

6.00 5.00 CS 4.50 3.00 2.00 1.50 0.50 9 87 6 -5 -4 -3 2 -1 0 1 23 4 5 6 7 8 9 10 11 12 Quantity